# Exact Factorization of Unitary Transformations with Spin-Adapted Generators

### v1

@@ Appendix B Seniority 4, intermediate triplet
The following linear combinations form the basis for the center ($Z(\mathcal{A}_{S4T})$)
$Z_{1}$ $=A_{5}+A_{21}+A_{29}+A_{35}+A_{41}+A_{99}+A_{101}-3A_{104}+A_{105}+A_{113}+A_{115}+A_{129}$
$Z_{2}$ $=A_{6}+A_{22}+A_{30}+A_{36}+A_{42}+A_{100}+A_{102}-3A_{103}+A_{106}+A_{114}+A_{116}+A_{130}$
$Z_{3}$ $=-A_{2}+A_{33}+A_{40}+A_{45}+A_{52}+A_{121}+3A_{124}-A_{127}-A_{128}-A_{133}-A_{134}+A_{135}$
$Z_{4}$ $=-A_{1}+A_{31}+A_{39}+A_{43}+A_{51}+A_{122}-A_{123}+3A_{125}-A_{126}-A_{131}-A_{132}+A_{136}$
$Z_{5}$ $=-A_{4}-A_{20}-A_{28}+A_{49}+A_{54}-A_{98}-3A_{108}+A_{109}+A_{112}+A_{118}+A_{120}+A_{137}$
$Z_{6}$ $=-A_{3}-A_{19}-A_{27}+A_{47}+A_{53}-A_{97}+A_{107}-3A_{110}+A_{111}+A_{117}+A_{119}+A_{138}.$
The central directions have the following explicit form
$Z_{1}$ $=\Big[\big(1-\hat{n}(i_{\beta})\big)\big(1-\hat{n}(j_{\beta})\big)\big(1-\hat{n}(a_{\beta})\big)\big(1-\hat{n}(b_{\beta})\big)+\hat{n}(i_{\beta})\hat{n}(j_{\beta})\hat{n}(a_{\beta})\hat{n}(b_{\beta})\Big]\,\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$Z_{2}$ $=\Big[\big(1-\hat{n}(i_{\alpha})\big)\big(1-\hat{n}(j_{\alpha})\big)\big(1-\hat{n}(a_{\alpha})\big)\big(1-\hat{n}(b_{\alpha})\big)+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\Big]\,\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$Z_{5}$ $=-\Big(\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})\big(1-\hat{n}(b_{\beta})\big)\big(1-\hat{n}(i_{\beta})\big)+\hat{n}(i_{\beta})\hat{n}(b_{\beta})\big(1-\hat{n}(j_{\alpha})\big)\big(1-\hat{n}(a_{\alpha})\big)\Big)\,\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$Z_{6}$ $=-\Big(\hat{n}(i_{\alpha})\hat{n}(b_{\alpha})\big(1-\hat{n}(a_{\beta})\big)\big(1-\hat{n}(j_{\beta})\big)+\hat{n}(j_{\beta})\hat{n}(a_{\beta})\big(1-\hat{n}(i_{\alpha})\big)\big(1-\hat{n}(b_{\alpha})\big)\Big)\,\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}.$
The semisimple part decomposes as a direct sum of pairwise commuting ideals, listed by bases in the $\{A_{i}\}$:
$I^{(3)}_{1}=\{\,A_{10},A_{22},A_{51}\,\}$
$I^{(3)}_{2}=\{\,A_{13},A_{29},A_{45}\,\}$
$I^{(3)}_{3}=\{\,A_{14},A_{30},A_{52}\,\}$
$I^{(3)}_{4}=\{\,A_{16},A_{36},A_{53}\,\}$
$I^{(3)}_{5}=\{\,A_{17},A_{41},A_{49}\,\}$
$I^{(3)}_{6}=\{\,A_{18},A_{42},A_{54}\,\}$
$I^{(3)}_{7}=\{\,A_{7},A_{19},A_{31}\,\}$
$I^{(6)}_{1}=\{\,A_{63},A_{69},A_{83},A_{101},A_{107},A_{123}\,\}$
$I^{(6)}_{2}=\{\,A_{64},A_{73},A_{84},A_{102},A_{111},A_{126}\,\}$
$I^{(12)}_{1}=\{\,A_{65},A_{66},A_{70},A_{72},A_{85},A_{86},A_{103},A_{104},A_{108},A_{110},A_{124},A_{125}\,\}$
$I^{(12)}_{2}=\{\,A_{23},A_{24},A_{32},A_{34},A_{55},A_{56},A_{59},A_{60},A_{97},A_{98},A_{121},A_{122}\,\}$
$I^{(15)}_{1}=\{\,A_{8},A_{20},A_{39},A_{67},A_{68},A_{71},A_{74},A_{93},A_{94},A_{105},A_{106},A_{109},A_{112},A_{131},A_{132}\,\}$
$I^{(15)}_{2}=\{\,A_{9},A_{21},A_{25},A_{26},A_{43},A_{44},A_{46},A_{57},A_{58},A_{61},A_{62},A_{99},A_{100},A_{135},A_{136}\,\}$
$I^{(15)}_{3}=\{\,A_{11},A_{27},A_{33},A_{75},A_{76},A_{79},A_{81},A_{87},A_{88},A_{113},A_{114},A_{117},A_{119},A_{127},A_{128}\,\}$
$I^{(15)}_{4}=\{\,A_{12},A_{28},A_{40},A_{77},A_{78},A_{80},A_{82},A_{95},A_{96},A_{115},A_{116},A_{118},A_{120},A_{133},A_{134}\,\}$
$I^{(15)}_{5}=\{\,A_{15},A_{35},A_{37},A_{38},A_{47},A_{48},A_{50},A_{89},A_{90},A_{91},A_{92},A_{129},A_{130},A_{137},A_{138}\,\}$
Here is the final direct sum of the algebra
$\mathcal{A}_{S4T}\;=\;\Big(\bigoplus_{i=1}^{7}I^{(3)}_{i}\;\oplus\;\bigoplus_{j=1}^{2}I^{(6)}_{j}\;\oplus\;\bigoplus_{k=1}^{2}I^{(12)}_{k}\;\oplus\;\bigoplus_{\ell=1}^{5}I^{(15)}_{\ell}\Big)\;\oplus\;Z(\mathcal{A}_{S4T}),$
with $Z(\mathcal{A}_{S4T})$ central and commuting with the semisimple part.

### v2

@@ Appendix B Seniority 4, intermediate triplet
The algebra $\mathcal{A}_{S4T}$ is reductive with
$\dim\mathcal{A}_{S4T}=84,\qquad\dim Z(\mathcal{A}_{S4T})=6,\qquad\dim[\mathcal{A}_{S4T},\mathcal{A}_{S4T}]=78.$
It decomposes as a direct sum of ideals
$\mathcal{A}_{S4T}\;=\;Z(\mathcal{A}_{S4T})\;\oplus\;\bigoplus_{r=1}^{26}\mathfrak{s}_{r},$
where each $\mathfrak{s}_{r}$ is a $3$-dimensional simple ideal (isomorphic to $\mathfrak{sl}_{2}$ over $\mathbb{C}$), and $\mathfrak{s}_{r},\mathfrak{s}_{r^{\prime}}]=0$ for $r\neq r^{\prime}$.
Center $Z(\mathcal{A}_{S4T})$
A basis $\{Z_{1},\dots,Z_{6}\}$ of the center is:
$Z_{1}$ $=2A_{3}+2A_{19}+A_{24}+2A_{27}+2A_{48}+A_{73}$
$Z_{2}$ $=2A_{4}+2A_{20}+A_{23}+2A_{28}+2A_{46}+A_{74}$
$Z_{3}$ $=2A_{5}+2A_{21}+A_{26}+2A_{29}-2A_{38}+A_{75}$
$Z_{4}$ $=2A_{6}+2A_{22}+A_{25}+2A_{30}-2A_{37}+A_{76}$
$Z_{5}$ $=-2A_{2}+A_{32}+2A_{33}+2A_{40}-2A_{42}+A_{81}$
$Z_{6}$ $=-2A_{1}+2A_{31}+A_{34}+2A_{39}-2A_{44}+A_{82}$
Semisimple part $\mathcal{A}_{S4T},\mathcal{A}_{S4T}]$
We list the $26$ commuting simple ideals $\mathfrak{s}_{r}$ (each of dimension $3$) via generators
$\mathfrak{s}_{r}=\mathrm{span}\{B^{(r)}_{1},B^{(r)}_{2},B^{(r)}_{3}\},\qquad B^{(r)}_{m}=\sum_{k}c^{(r,m)}_{k}A_{k}.$
Ideal $\mathfrak{s}_{1}$ (dim 33).
$B^{(1)}_{1}$ $=-A_{32}-A_{34}+2A_{67}+2A_{68}+A_{81}+A_{82}$
$B^{(1)}_{2}$ $=-A_{25}-A_{26}-A_{49}-A_{50}-A_{53}-A_{54}+A_{75}+A_{76}$
$B^{(1)}_{3}$ $=-A_{23}-A_{24}+A_{51}+A_{52}+A_{55}+A_{56}+A_{73}+A_{74}$
Ideal $\mathfrak{s}_{2}$ (dim 33).
$B^{(2)}_{1}$ $=-A_{32}-A_{34}-2A_{67}-2A_{68}+A_{81}+A_{82}$
$B^{(2)}_{2}$ $=-A_{25}-A_{26}+A_{49}+A_{50}+A_{53}+A_{54}+A_{75}+A_{76}$
$B^{(2)}_{3}$ $=-A_{23}-A_{24}-A_{51}-A_{52}-A_{55}-A_{56}+A_{73}+A_{74}$
Ideal $\mathfrak{s}_{3}$ (dim 33).
$B^{(3)}_{1}$ $=A_{32}-A_{34}+2A_{67}-2A_{68}-A_{81}+A_{82}$
$B^{(3)}_{2}$ $=-A_{25}+A_{26}+A_{49}-A_{50}-A_{53}+A_{54}-A_{75}+A_{76}$
$B^{(3)}_{3}$ $=-A_{23}+A_{24}-A_{51}+A_{52}+A_{55}-A_{56}-A_{73}+A_{74}$
Ideal $\mathfrak{s}_{4}$ (dim 33).
$B^{(4)}_{1}$ $=A_{32}-A_{34}-2A_{67}+2A_{68}-A_{81}+A_{82}$
$B^{(4)}_{2}$ $=-A_{25}+A_{26}-A_{49}+A_{50}+A_{53}-A_{54}-A_{75}+A_{76}$
$B^{(4)}_{3}$ $=-A_{23}+A_{24}+A_{51}-A_{52}-A_{55}+A_{56}-A_{73}+A_{74}$
Ideal $\mathfrak{s}_{5}$ (dim 33).
$B^{(5)}_{1}$ $=-\frac{1}{2}A_{34}-A_{65}+A_{68}+\frac{1}{2}A_{82}+A_{83}$
$B^{(5)}_{2}$ $=-\frac{1}{2}A_{24}+\frac{1}{2}A_{52}+\frac{1}{2}A_{55}-A_{57}+\frac{1}{2}A_{73}+A_{79}$
$B^{(5)}_{3}$ $=\frac{1}{2}A_{26}+\frac{1}{2}A_{50}+\frac{1}{2}A_{53}+A_{59}-\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{6}$ (dim 33).
$B^{(6)}_{1}$ $=-\frac{1}{2}A_{34}+A_{65}-A_{68}+\frac{1}{2}A_{82}+A_{83}$
$B^{(6)}_{2}$ $=-\frac{1}{2}A_{24}-\frac{1}{2}A_{52}-\frac{1}{2}A_{55}+A_{57}+\frac{1}{2}A_{73}+A_{79}$
$B^{(6)}_{3}$ $=\frac{1}{2}A_{26}-\frac{1}{2}A_{50}-\frac{1}{2}A_{53}-A_{59}-\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{7}$ (dim 33).
$B^{(7)}_{1}$ $=-A_{31}-\frac{1}{2}A_{34}-A_{66}+A_{67}-\frac{1}{2}A_{82}+A_{83}$
$B^{(7)}_{2}$ $=A_{7}-A_{19}-\frac{1}{2}A_{24}-\frac{1}{2}A_{25}-\frac{1}{2}A_{50}+\frac{1}{2}A_{51}+\frac{1}{2}A_{53}+\frac{1}{2}A_{56}-A_{58}-A_{59}-\frac{1}{2}A_{73}+\frac{1}{2}A_{76}-A_{78}+A_{79}$
$B^{(7)}_{3}$ $=-A_{7}+\frac{1}{2}A_{25}+\frac{1}{2}A_{50}-\frac{1}{2}A_{53}+A_{59}-\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{8}$ (dim 33).
$B^{(8)}_{1}$ $=-A_{31}-\frac{1}{2}A_{34}+A_{66}-A_{67}-\frac{1}{2}A_{82}+A_{83}$
$B^{(8)}_{2}$ $=-A_{19}-\frac{1}{2}A_{24}-\frac{1}{2}A_{51}-\frac{1}{2}A_{56}+A_{58}-\frac{1}{2}A_{73}+A_{79}$
$B^{(8)}_{3}$ $=A_{7}+\frac{1}{2}A_{25}-\frac{1}{2}A_{50}+\frac{1}{2}A_{53}-A_{59}-\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{9}$ (dim 33).
$B^{(9)}_{1}$ $=A_{32}+A_{34}+A_{81}+A_{82}$
$B^{(9)}_{2}$ $=A_{23}-A_{24}-A_{73}+A_{74}$
$B^{(9)}_{3}$ $=-A_{49}+A_{50}-A_{53}+A_{54}$
Ideal $\mathfrak{s}_{10}$ (dim 33).
$B^{(10)}_{1}$ $=-A_{32}+A_{34}-A_{81}+A_{82}$
$B^{(10)}_{2}$ $=A_{23}+A_{24}+A_{73}+A_{74}$
$B^{(10)}_{3}$ $=-A_{49}-A_{50}+A_{53}+A_{54}$
Ideal $\mathfrak{s}_{11}$ (dim 33).
$B^{(11)}_{1}$ $=\frac{1}{2}A_{34}-A_{41}-A_{44}+A_{67}-A_{71}-\frac{1}{2}A_{82}+A_{83}$
$B^{(11)}_{2}$ $=-A_{9}-\frac{1}{2}A_{23}+\frac{1}{2}A_{52}-\frac{1}{2}A_{55}-A_{57}+\frac{1}{2}A_{74}+A_{80}$
$B^{(11)}_{3}$ $=A_{21}+\frac{1}{2}A_{26}+\frac{1}{2}A_{49}+\frac{1}{2}A_{54}+A_{60}+\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{12}$ (dim 33).
$B^{(12)}_{1}$ $=\frac{1}{2}A_{34}-A_{41}-A_{44}-A_{67}+A_{71}-\frac{1}{2}A_{82}+A_{83}$
$B^{(12)}_{2}$ $=A_{9}-\frac{1}{2}A_{23}-\frac{1}{2}A_{52}+\frac{1}{2}A_{55}+A_{57}+\frac{1}{2}A_{74}+A_{80}$
$B^{(12)}_{3}$ $=A_{21}+\frac{1}{2}A_{26}-\frac{1}{2}A_{49}-\frac{1}{2}A_{54}-A_{60}+\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{13}$ (dim 33).
$B^{(13)}_{1}$ $=\frac{1}{2}A_{34}+A_{39}-A_{41}-A_{44}+A_{68}-A_{72}+\frac{1}{2}A_{82}+A_{83}$
$B^{(13)}_{2}$ $=-A_{10}-A_{20}-\frac{1}{2}A_{23}+\frac{1}{2}A_{51}-\frac{1}{2}A_{56}-A_{58}-\frac{1}{2}A_{74}+A_{80}$
$B^{(13)}_{3}$ $=-A_{8}+A_{22}+\frac{1}{2}A_{25}+\frac{1}{2}A_{49}-\frac{1}{2}A_{54}+A_{60}+\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{14}$ (dim 33).
$B^{(14)}_{1}$ $=\frac{1}{2}A_{34}+A_{39}-A_{41}-A_{44}-A_{68}+A_{72}+\frac{1}{2}A_{82}+A_{83}$
$B^{(14)}_{2}$ $=A_{10}-A_{20}-\frac{1}{2}A_{23}-\frac{1}{2}A_{51}+\frac{1}{2}A_{56}+A_{58}-\frac{1}{2}A_{74}+A_{80}$
$B^{(14)}_{3}$ $=A_{8}+A_{22}+\frac{1}{2}A_{25}-\frac{1}{2}A_{49}+\frac{1}{2}A_{54}-A_{60}+\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{15}$ (dim 33).
$B^{(15)}_{1}$ $=-A_{16}-A_{17}+\frac{1}{2}A_{32}-A_{42}-A_{43}-A_{66}+A_{67}-A_{70}-A_{71}-\frac{1}{2}A_{81}+A_{84}$
$B^{(15)}_{2}$ $=-A_{13}+\frac{1}{2}A_{23}+\frac{1}{2}A_{26}+A_{29}-A_{35}-A_{38}+A_{46}+A_{47}-\frac{1}{2}A_{50}-\frac{1}{2}A_{51}-\frac{1}{2}A_{53}+\frac{1}{2}A_{56}-A_{61}-A_{64}-\frac{1}{2}A_{74}+\frac{1}{2}A_{75}-A_{77}+A_{80}$
$B^{(15)}_{3}$ $=-\frac{1}{2}A_{26}-A_{29}+A_{35}+A_{38}+\frac{1}{2}A_{50}+\frac{1}{2}A_{53}+A_{64}-\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{16}$ (dim 33).
$B^{(16)}_{1}$ $=A_{16}+A_{17}+\frac{1}{2}A_{32}-A_{42}-A_{43}+A_{66}-A_{67}+A_{70}+A_{71}-\frac{1}{2}A_{81}+A_{84}$
$B^{(16)}_{2}$ $=A_{13}+\frac{1}{2}A_{23}+A_{46}+A_{47}+\frac{1}{2}A_{51}-\frac{1}{2}A_{56}+A_{61}-\frac{1}{2}A_{74}+A_{80}$
$B^{(16)}_{3}$ $=-\frac{1}{2}A_{26}-A_{29}+A_{35}+A_{38}-\frac{1}{2}A_{50}-\frac{1}{2}A_{53}-A_{64}-\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{17}$ (dim 33).
$B^{(17)}_{1}$ $=-A_{15}-A_{18}+\frac{1}{2}A_{32}+A_{40}-A_{42}-A_{43}-A_{65}+A_{68}-A_{69}-A_{72}+\frac{1}{2}A_{81}+A_{84}$
$B^{(17)}_{2}$ $=-A_{14}+\frac{1}{2}A_{23}+A_{28}+A_{46}+A_{47}-\frac{1}{2}A_{52}+\frac{1}{2}A_{55}-A_{62}+\frac{1}{2}A_{74}+A_{80}$
$B^{(17)}_{3}$ $=-A_{12}-\frac{1}{2}A_{25}-A_{30}+A_{36}+A_{37}-\frac{1}{2}A_{50}+\frac{1}{2}A_{53}+A_{64}-\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{18}$ (dim 33).
$B^{(18)}_{1}$ $=A_{15}+A_{18}+\frac{1}{2}A_{32}+A_{40}-A_{42}-A_{43}+A_{65}-A_{68}+A_{69}+A_{72}+\frac{1}{2}A_{81}+A_{84}$
$B^{(18)}_{2}$ $=A_{14}+\frac{1}{2}A_{23}+A_{28}+A_{46}+A_{47}+\frac{1}{2}A_{52}-\frac{1}{2}A_{55}+A_{62}+\frac{1}{2}A_{74}+A_{80}$
$B^{(18)}_{3}$ $=A_{12}-\frac{1}{2}A_{25}-A_{30}+A_{36}+A_{37}+\frac{1}{2}A_{50}-\frac{1}{2}A_{53}-A_{64}-\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{19}$ (dim 33).
$B^{(19)}_{1}$ $=-A_{25}+A_{26}-A_{32}-A_{34}+2A_{42}+2A_{44}+A_{75}-A_{76}+A_{81}+A_{82}$
$B^{(19)}_{2}$ $=A_{25}-A_{26}-A_{75}+A_{76}$
$B^{(19)}_{3}$ $=-A_{51}+A_{52}-A_{55}+A_{56}$
Ideal $\mathfrak{s}_{20}$ (dim 33).
$B^{(20)}_{1}$ $=A_{32}-A_{34}-2A_{42}+2A_{44}-A_{81}+A_{82}$
$B^{(20)}_{2}$ $=A_{25}+A_{26}+A_{75}+A_{76}$
$B^{(20)}_{3}$ $=-A_{51}-A_{52}+A_{55}+A_{56}$
Ideal $\mathfrak{s}_{21}$ (dim 33).
$B^{(21)}_{1}$ $=-A_{25}+A_{26}+2A_{37}-2A_{38}-A_{75}+A_{76}$
$B^{(21)}_{2}$ $=-A_{23}-A_{24}-2A_{46}-2A_{48}+A_{73}+A_{74}$
$B^{(21)}_{3}$ $=-A_{15}+A_{16}-A_{65}+A_{66}-A_{67}+A_{68}-A_{69}+A_{70}$
Ideal $\mathfrak{s}_{22}$ (dim 33).
$B^{(22)}_{1}$ $=-A_{25}-A_{26}+2A_{37}+2A_{38}+A_{75}+A_{76}$
$B^{(22)}_{2}$ $=-A_{23}+A_{24}-2A_{46}+2A_{48}-A_{73}+A_{74}$
$B^{(22)}_{3}$ $=A_{15}+A_{16}+A_{65}+A_{66}-A_{67}-A_{68}+A_{69}+A_{70}$
Ideal $\mathfrak{s}_{23}$ (dim 33).
$B^{(23)}_{1}$ $=-\frac{1}{2}A_{32}+A_{68}-A_{69}+\frac{1}{2}A_{81}+A_{84}$
$B^{(23)}_{2}$ $=\frac{1}{2}A_{24}+A_{45}+A_{48}-\frac{1}{2}A_{51}-\frac{1}{2}A_{56}+A_{61}-\frac{1}{2}A_{73}+A_{79}$
$B^{(23)}_{3}$ $=-\frac{1}{2}A_{26}+A_{35}+A_{38}-\frac{1}{2}A_{49}-\frac{1}{2}A_{54}-A_{63}+\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{24}$ (dim 33).
$B^{(24)}_{1}$ $=-\frac{1}{2}A_{32}-A_{68}+A_{69}+\frac{1}{2}A_{81}+A_{84}$
$B^{(24)}_{2}$ $=\frac{1}{2}A_{24}+A_{45}+A_{48}+\frac{1}{2}A_{51}+\frac{1}{2}A_{56}-A_{61}-\frac{1}{2}A_{73}+A_{79}$
$B^{(24)}_{3}$ $=-\frac{1}{2}A_{26}+A_{35}+A_{38}+\frac{1}{2}A_{49}+\frac{1}{2}A_{54}+A_{63}+\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{25}$ (dim 33).
$B^{(25)}_{1}$ $=-\frac{1}{2}A_{32}-A_{33}+A_{67}-A_{70}-\frac{1}{2}A_{81}+A_{84}$
$B^{(25)}_{2}$ $=\frac{1}{2}A_{24}+A_{27}+A_{45}+A_{48}-\frac{1}{2}A_{52}-\frac{1}{2}A_{55}+A_{62}+\frac{1}{2}A_{73}+A_{79}$
$B^{(25)}_{3}$ $=A_{11}-\frac{1}{2}A_{25}+A_{36}+A_{37}+\frac{1}{2}A_{49}-\frac{1}{2}A_{54}-A_{63}+\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{26}$ (dim 33).
$B^{(26)}_{1}$ $=-\frac{1}{2}A_{32}-A_{33}-A_{67}+A_{70}-\frac{1}{2}A_{81}+A_{84}$
$B^{(26)}_{2}$ $=\frac{1}{2}A_{24}+A_{27}+A_{45}+A_{48}+\frac{1}{2}A_{52}+\frac{1}{2}A_{55}-A_{62}+\frac{1}{2}A_{73}+A_{79}$
$B^{(26)}_{3}$ $=-A_{11}-\frac{1}{2}A_{25}+A_{36}+A_{37}-\frac{1}{2}A_{49}+\frac{1}{2}A_{54}+A_{63}+\frac{1}{2}A_{76}+A_{78}$